OpenStudy (anonymous):

f(x)=sin(^2)x-3sinxcosx+2cos(^2)x g(x)=f(x)-f(Pi/2-x) Find g(Pi/8)

OpenStudy (anonymous):

f(Pi/2-x) change sines to cosine

OpenStudy (anonymous):

f(Pi/2 -x)= Cos^2(x)- 3 sin(x) cos(x)+ 2 Sin^2(x) so it turn out to be the same

OpenStudy (anonymous):

f(x)- f(Pi/2 -x )=0 since they are the same

OpenStudy (anonymous):

Thanks!